OpenStudy (anonymous):

convert 0.00311m to nanometers

OpenStudy (anonymous):

0.00311m X 10^9=3110000 um

OpenStudy (aaronq):

^ thats correct. except for his units, they should read "nm" not "um".

OpenStudy (anonymous):

haha thats right my bad

OpenStudy (aaronq):

haha good job, regardless

OpenStudy (unklerhaukus):

\[1[\text{nm}]=10^{-9}[\text m]\quad\iff\quad10^{9}\frac{[\text{nm}]}{[\text m]}=1\] \[ 0.00311[\text m]\]\[=0.00311[\text m]\times1\]\[=0.00311[\text m]\times10^{9}\frac{[\text{nm}]}{[\text m]}\]\[=\left(0.00311\times10^9\right)\cancel{[\text m]}\frac{[\text{nm}]}{\cancel{[\text m]}}\]\[=0.00311\times10^9[\text{nm}]\]\[=3.11\times10^{-3}\times10^9[\text{nm}]\]\[=3.11\times10^6[\text{nm}]\]

OpenStudy (preetha):

Wow! Unkle, that is one fine formatted answer!

OpenStudy (unklerhaukus):

\[\color{teal}{\ddot\smile}\]

Parth (parthkohli):

\[\Huge \stackrel{\rm Unkle}{ \bf is}{the}{ ~\\\LaTeX~\text{god!}}\]

Parth (parthkohli):

@raeganbanks Hey, welcome. This question is one of the general reference ones which best fits in Mathematics. :-)

OpenStudy (unklerhaukus):

\[\color{brown}{\large\ddot\eta}\]

Parth (parthkohli):

\[\ddot\smile \]

OpenStudy (unklerhaukus):

\[\stackrel {\therefore}\smile\]

Parth (parthkohli):

\[\stackrel{\vee }{\ddot \smile }\]

Parth (parthkohli):

\[\ddot \omega\]

Parth (parthkohli):

\[\stackrel{\Large \varpi}{\smile}\]

OpenStudy (unklerhaukus):

\[\color{cyan}{\langle\check{\color{gray}{\ddot \lor}}\rangle}\]

Parth (parthkohli):

\[\boxed{\stackrel{\vee}{\ddot \smile }}\]

OpenStudy (unklerhaukus):

\[\boxed{\color{teal}{\langle\!\color{brown}{\langle}\check{\color{grey}{\ddot\lor}}\color{brown}{\rangle}\!\rangle}}\]